OpenStudy (anonymous):

Frank needs to convert 24 centimeters per second to meters per minute. Which conversion factors should she use? A. B. C. D.

OpenStudy (gorv):

\[\frac{ cm }{ \sec }=\frac{ 10^{-2} m }{ \frac{ 1 }{ 60 } milute }\]

OpenStudy (gorv):

\[that is 60*10^{-2}=0.6\]

OpenStudy (gorv):

@unicornsoup